A sinusoidal wave travelling in the positive x direction has an amllplitude of 15.0 cm, a wavelength of 40.0 cm, and a frequency pf 8.00 Hz. The vertical position of an element of the medium at t=0 and x =0 is also 15.0 cm.

A. Find the wave number k, period T, angular frequency w, and speed v of the wave

B. Write a general expression for the wave function


1
Expert's answer
2021-01-29T20:01:39-0500

Let's denote: "A = 15cm =0.15m" is the amplitude, "\lambda = 40cm = 0.4m" is the wavelength, "f = 8Hz" is the frequency.

The wave number is given as:


"k = \dfrac{2\pi}{\lambda} = \dfrac{2\pi}{0.4} = 5\pi\space m^{-1}"

The angular frequency is:


"\omega = 2\pi f = 16\pi \space rad/s"

The speed is:


"v = \lambda f = 3.2m/s"

The general expression for the wave is:


"y(x,t) = A\cos(kx - \omega t)\\
y(x,t) = 0.15\cos(5\pi x - 16\pi t)\space [m]"

Answer. A) "k = 5\pi \space m^{-1}, \omega = 16\pi \space rad/s, v = 3.2m/s", B) "y(x,t) = 0.15\cos(5\pi x - 16\pi t)\space [m]".